# Vector Quantization using Gaussian Variational Autoencoder

### v1

@@ D.10 Complexity
As with FSQ and BSQ (mentzer2023finite; zhao2024image), our codebook can be generated on the fly by maintaining the same random number generator seed on both the encoder and decoder sides. Therefore, our GQ model has the same parameter size as the vanilla Gaussian VAE. In Table 22, we compare the encoding and decoding frames per second (FPS) of the Gaussian VAE and GQ. We use $256\times 256$ images with a batch size of $1$, and we report the wall clock time, meaning that the time required for loading data is included. The results show that the encoding FPS of GQ (implemented in PyTorch) is 12 on an H100 GPU, which is considerably slower than the 104 FPS achieved by the Gaussian VAE. On the other hand, GQ does not incur any decoding overhead.
@@
To reduce the computational complexity of GQ, we implement GQ using a tailored CUDA kernel. Specifically, we follow the approach of Vonderfecht2025LossyCW, with a key difference: we maintain the codebook, as our bottleneck is not codebook instantiation. Additionally, we avoid the creation of large buffer vectors by performing the summation over $m$ within the CUDA kernel instead of in PyTorch. With this approach, we achieve an encoding FPS of approximately 80, with negligible overhead compared to the Gaussian VAE. A detailed comparison between the PyTorch implementation and the CUDA implementation of GQ is provided below as GQ_torch and GQ_CUDA, respectively.
1def GQ_torch(mu, sigma, codebook, m, bs, K):
2 # mu.shape = (bs, m)
3 # sigma.shape = (bs, m)
4 # codebook.shape = (K, m)
5
6 # This step create (bs, m, K) vector, which is the performance bottleneck
7 dist_m =((mu[:,None] - codebook[None])/sigma[:,None])**2
8 dist = torch.sum(dist_m, dim=1) # sum over m dimension
9 indices = torch.argmin(dist, dim=1) # argmax over K dimension
@@
10 zhat = torch.index_select(codebook, 0, indices) # select quantized results
11 return indices, zhat
1def GQ_CUDA(mu, sigma, codebook, m, bs, K):
2 dist = torch.zeros([bs, K])
@@
3 # need an extension wrapping and register the kernel into operator, we omit it in paper
4 # see code appendix for details
5 GQ_Kernel<<<bs * K / 256,256>>>(mu, sigma, codebook, dist, m, bs, K)
6
7 indices = torch.argmin(dist, dim=1) # argmax over K dimension
@@ 5 GQ_Kernel<<<bs * K / 256,256>>>(mu, sigma, codebook, dist, m, bs, K)
8 zhat = torch.index_select(codebook, 0, indices) # select quantized results
9 return indices, zhat
10
11__global__ void GQ_Kernel(
12 const float* mu,
13 const float* sigma,
14 const float* codebook,
15 float* dist,
16 int64_t m,
17 int64_t bs,
18 int64_t K
19) {
20 int idx = blockIdx.x * blockDim.x + threadIdx.x;
21 if (idx >= K * bs) return;
22 int bi = idx / K;
23 int ni = idx % K;
24 float a = 0.0f;
25 for (int i = 0; i < m; i++) {
26 float b = (codebook[ni * m + i] - mu[bi * dim + i]) / sigma[bi * dim + i];
27 a += b * b;
28 }
29 dist[idx] = a;
30 return;
31}
D.11 Asymptotic Complexity
@@
It is noteworthy that GQ without grouping quantization, such as PT or TR, or GQ with a group size of $m=1$, is asymptotically faster than reverse channel coding methods. This is because, for $m=1$, the GQ target in Eq.3 reduces to a quadratic form. In this case, it suffices to sort the scalar codebook $c_{1:K}$ in advance. Despite the sorting takes $\Omega(D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)))$, the sorting is only need to be done once and can be amortized across dimension and dataset. Subsequently, the minimization in Eq.3 can be performed in $O(D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ time using binary search. The details is shown in Algorithm 2.
@@
On the other hand, most reverse channel coding methods require $O(2^{D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1))})$ computational complexity (havasi2018minimal; Flamich2020CompressingIB; Theis2021AlgorithmsFT). A coding (Flamich2022FastRE) can achieve $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ encoding complexity, albeit at the cost of increased decoding complexity.
@@ E.1 Additional Qualitative Results and Failure Cases
In Figure 5, we present additional qualitative results showing that GQ achieves superior visual quality. However, we also note that none of the approaches is successful in reconstructing the plate of the residential vehicle. The text content remains challenging for low bitrate VQ-VAEs.
1input $c_{1:K}(\textrm{sorted, such that }c_{j}\leq c_{j+1}),\mu_{i}$
@@
3for $j=1$ to $K$ do
[…]
6return $c_{j^{*}},j^{*}$
Algorithm 1 GQ (argmax)
1input $c_{1:K}(\textrm{sorted, such that }c_{j}\leq c_{j+1}),\mu_{i}$
2l=1,r=Kl=1,r=K
3while $l+1<r$ do
4   $m=(l+r)//2$
5   if $c_{m}<\mu_{i}$ do
6    $l=m$
7   else
8    $r=m$
[…]
10   return $c_{l},l$
11else
12   return $c_{r},r$
@@ 8    $r=m$
Figure 5: Qualitative results on ImageNet dataset and 0.25 bpp. None of those approaches correctly reconstruct the plate number.

### v2

@@ D.6 Complexity
As with FSQ and BSQ (Mentzer et al., 2023; Zhao et al., 2024), our codebook can be generated on the fly by maintaining the same random number generator seed on both the encoder and decoder sides. Therefore, our GQ model has the same parameter size as the vanilla Gaussian VAE. In Table 20, we compare the encoding and decoding frames per second (FPS) of the Gaussian VAE and GQ. We use $256\times 256$ images with a batch size of $1$, and we report the wall clock time, meaning that the time required for loading data is included. The results show that the encoding FPS of GQ (implemented in PyTorch) is 12 on an H100 GPU, which is considerably slower than the 104 FPS achieved by the Gaussian VAE. On the other hand, GQ does not incur any decoding overhead.
@@
To reduce the computational complexity of GQ, we implement GQ using a tailored CUDA kernel. Specifically, we follow the approach of Vonderfecht and Liu (2025), with a key difference: we maintain the codebook, as our bottleneck is not codebook instantiation. Additionally, we avoid the creation of large buffer vectors by performing the summation over $m$ within the CUDA kernel instead of in PyTorch. With this approach, we achieve an encoding FPS of approximately 80, with negligible overhead compared to the Gaussian VAE. A detailed comparison between the PyTorch implementation and the CUDA implementation of GQ is provided below as GQ_torch and GQ_CUDA, respectively.
def GQ_torch(mu, sigma, codebook, m, bs, K):
# mu.shape = (bs, m)
# sigma.shape = (bs, m)
# codebook.shape = (K, m)
# This step create (bs, m, K) vector, which is the performance bottleneck
dist_m =((mu[:,None] - codebook[None])/sigma[:,None])**2
dist = torch.sum(dist_m, dim=1) # sum over m dimension
indices = torch.argmin(dist, dim=1) # argmax over K dimension
@@
zhat = torch.index_select(codebook, 0, indices) # select quantized results
return indices, zhat
def GQ_CUDA(mu, sigma, codebook, m, bs, K):
dist = torch.zeros([bs, K])
@@
# need an extension wrapping and register the kernel into operator, we omit it in paper
# see code appendix for details
GQ_Kernel<<<bs * K / 256,256>>>(mu, sigma, codebook, dist, m, bs, K)
indices = torch.argmin(dist, dim=1) # argmax over K dimension
@@
zhat = torch.index_select(codebook, 0, indices) # select quantized results
return indices, zhat
__global__ void GQ_Kernel(
const float* mu,
const float* sigma,
const float* codebook,
float* dist,
int64_t m,
int64_t bs,
int64_t K
int idx = blockIdx.x * blockDim.x + threadIdx.x;
if (idx >= K * bs) return;
int bi = idx / K;
int ni = idx % K;
float a = 0.0f;
for (int i = 0; i < m; i++) {
float b = (codebook[ni * m + i] - mu[bi * dim + i]) / sigma[bi * dim + i];
a += b * b;
dist[idx] = a;
return;
D.7 Asymptotic Complexity
@@
It is noteworthy that GQ with codebook dimension $m=1$, is asymptotically faster than reverse channel coding methods. This is because, for $m=1$, the GQ target in Eq.3 reduces to a quadratic form. In this case, it suffices to sort the scalar codebook $c_{1:K}$ in advance. Despite the sorting takes $\Omega(D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)))$, the sorting is only need to be done once and can be amortized across dimension and dataset. Subsequently, the minimization in Eq.3 can be performed in $O(D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ time using binary search. The details is shown in Algorithm 2.
@@
On the other hand, most reverse channel coding methods require $O(2^{D_{\mathrm{KL}}(q(Z_{i}|X)||\mathcal{N}(0,1))})$ computational complexity (Havasi et al., 2018a; Flamich et al., 2020; Theis and Yosri, 2021). A coding (Flamich et al., 2022) can achieve $O(D_{\infty}(q(Z_{i}|X)||\mathcal{N}(0,1)))$ encoding complexity, albeit at the cost of increased decoding complexity.
@@ E.1 Additional Qualitative Results and Failure Cases
In Figure 4, we present additional qualitative results showing that GQ achieves superior visual quality. However, we also note that none of the approaches is successful in reconstructing the license of the residential vehicle. The text content remains challenging for low bitrate VQ-VAEs.
Algorithm 1 GQ (Argmax)
0: Codebook $c_{1:K}$ (sorted, $c_{j}\leq c_{j+1}$), $\mu_{i}$
0: Quantized value $\hat{z}_{i}$, index $j^{*}$
@@
2:for $j=1$ to $K$ do
3:  if $T^{*}\leq\|c_{j}-\mu_{i}\|$ then
4:   $T^{*}\leftarrow\|c_{j}-\mu_{i}\|$
5:   $j^{*}\leftarrow j$
6:  end if
7:end for
8:return $c_{j^{*}},j^{*}$
@@
0: Codebook $c_{1:K}$ (sorted, $c_{j}\leq c_{j+1}$), $\mu_{i}$
0: Quantized value $\hat{z}_{i}$, index $l$ or $r$
$1l\leftarrow$ 1, $r\leftarrow K$
2:while $l+1<r$ do
3:  $m\leftarrow\lfloor(l+r)/2\rfloor$
4:  if $c_{m}<\mu_{i}$ then
5:   $l\leftarrow m$
6:  else
7:   $r\leftarrow m$
8:  end if
9:end while
10:if $\|c_{l}-\mu_{i}\|<\|c_{r}-\mu_{i}\|$ then
11:  return $c_{l},l$
12:else
13:  return $c_{r},r$
14:end if
Figure 4: Qualitative results on ImageNet dataset and 0.25 bpp. None of these approaches correctly reconstructs the license plate.